INFRARED SPECTRUM OF THE ν_{1} FUNDAMENTAL BAND OF H_{3}O^{+}

\underline{J. TANG} and T. OKA, Department of Chemistry, The University of Chicago, Chicago, IL 60637.

Infrared spectrum of H_{3}O^{+} has been observed by a difference frequency laser spectrometer with positive column discharges of H_{2}/O_{2} gas mixtures, which gave a spectrum remarkably simpler than the previous one with H_{2}/O_{2}/He discharges.^{a} After extending the previous assignment of the ν_{3} fundamental bands to higher J, K transitions, vibration-rotation structures due to the ν_{1}^{+} ← 0^{−} and ν_{1}^{−} ← 0^{+} bands of H_{3}O^{+} were identified with the band origins of 3389.66 cm^{−1} and 3491.17 cm^{−1}, respectively, in the region of the strong ν_{3}^{±} ← 0^{±} bands. Molecular constants for the ν_{1}^{±} ← 0^{∓} bands were obtained by the least-squares fittings for the observed frequencies. Coriolis interactions between the ν_{1} and ν_{3} vibrations have been considered to explain some large deviations of the observed transitions from the calculated frequencies in the ν_{1} and ν_{3} bands.

aW. C. Ho, C. J. Pursell, and T. Oka, J. Mol. Spectrosc. \underline{149}, 530 (1991).